# Know Your Limits: A Survey of Abstention in Large Language Models

### v2

@@ Abstract
Abstention, the refusal of large language models (LLMs) to provide an answer, is increasingly recognized for its potential to mitigate hallucinations and enhance safety in LLM systems. In this survey, we introduce a framework to examine abstention from three perspectives: the query, the model, and human values. We organize the literature on abstention methods, benchmarks, and evaluation metrics using this framework, and discuss merits and limitations of prior work. We further identify and motivate areas for future work, centered around whether abstention can be achieved as a meta-capability that transcends specific tasks or domains, while still providing opportunities to optimize abstention abilities based on context.
@@ 1 Introduction
Large language models (LLMs) have demonstrated generalization capabilities across NLP tasks such as question answering (QA) (Wei et al., 2022; Chowdhery et al., 2022), abstractive summarization (Zhang et al., 2023a), and dialogue generation (Yi et al., 2024). But these models are also unreliable, having a tendency to “hallucinate” false information in their responses Ji et al. (2023b), generate overly certain or authoritative responses Zhou et al. (2024b), answer with incomplete information Zhou et al. (2023b), or produce harmful or dangerous responses Anwar et al. (2024). In these situations, the model should ideally abstain: to refuse to answer in the face of uncertainty (Wen et al., 2024; Feng et al., 2024b; Yang et al., 2023).
@@
Current methods to encourage abstention typically rely on calibration techniques, including linguistic calibration (Mielke et al., 2022; Huang et al., 2024b), which aim to accurately and consistently estimate a model’s confidence in its response, then arrange for the model to abstain if the confidence score for a given response falls below some threshold threshold (Varshney et al., 2022; Xiao et al., 2022; Desai and Durrett, 2020). But questions of human values and the answerability of the query itself are difficult to model in terms of model confidence Yang et al. (2023).
While prior work demonstrates the potential of abstention in enhancing model safety and reliability Varshney et al. (2023); Wang et al. (2024c); Zhang et al. (2024a), the study of abstention has also been constrained to specific QA tasks. This task-specific approach limits the broader applicability of abstention strategies across the diverse range of scenarios encountered by general-purpose chatbots engaging in open-domain interactions. Recognizing these limitations, this paper aims to broaden the understanding of abstention by studying it within a general chatbot system framework. Our goals are to encourage new methods to achieve abstention and adapt abstention mechanisms to suit a wide array of tasks, thereby enhancing the overall robustness and trustworthiness of AI interactions.
@@
  • We introduce a framework to analyze abstention capabilities from three perspectives—the query, the model, and human values. This framework organizes existing research and provides a foundation for developing adaptable and robust abstention mechanisms for AI systems (§2).
  • We conduct a detailed survey of existing abstention methods (§3) as well as evaluation benchmarks and metrics (§4), aiding researchers in selecting appropriate strategies and identifying areas for further research to advance the field.
  • We discuss other considerations and underexplored aspects (§5) of abstention, highlighting pitfalls and promising future directions. We encourage researchers to develop more robust model abstention mechanisms and demonstrate their effectiveness in real-world applications.
@@ 2 Abstention in LLMs
We define abstention as the refusal to answer a query. Abstention includes a spectrum of behaviors ranging from partial to full abstention Röttger et al. (2024a); for example, expressing uncertainty, providing conflicting conclusions, or refusing to respond due to potential harm are all forms of abstention. When a model fully abstains, it may respond “I don’t know” or refuse to answer in another way. Whereas partial abstention may involve both answering and abstention, such as self-contradictory responses, e.g., “I can’t answer the question, but I suppose the answer might be…” We do not consider ignoring and/or reframing the question as abstention; but rather as failure modes of LLMs in following instructions (Röttger et al., 2024a; Varshney et al., 2023).
For the abstention expression—the words a model uses to convey that it has abstained—we adopt five major types of expressions from prior work (Varshney et al., 2023; Wang et al., 2024c), indicating that the model (i) cannot assist; (ii) refutes the query; (iii) provides multiple perspectives without expressing preference; (iv) perceives risk associated with the query and answers cautiously with a disclaimer; and (v) refuses to offer concrete answers due to lack of knowledge or certainty. Expressions can be identified through heuristic rules and key word matching (Zou et al., 2023; Wen et al., 2024; Yang et al., 2023), and through model-based or human-based evaluation (§4).
@@ 2.1 Abstention Framework
  • The model knowledge perspective examines the capabilities of the AI model itself, including its design, training, and inherent biases (Ahdritz et al., 2024; Kim and Thorne, 2024; Hestness et al., 2017; Hoffmann et al., 2022; Kaplan et al., 2020). For any given query, the system should abstain if the model is insufficiently confident about the correctness of output or has a high probability of returning an incorrect output.
@@ 2.2 Problem Formulation
Consider an LLM $f:\mathcal{X}\rightarrow\mathcal{Y}$. When given a prompt $\mathbf{x}\in\mathcal{X}$, $f$ generates a response $\mathbf{y}\in\mathcal{Y}$. We model refusal to answer (abstention) as a function $r:\mathcal{X,Y}\rightarrow\{0,1\}$ where $r(\mathbf{x},\mathbf{y})=1$ indicates the system will abstain from answering and $r(\mathbf{x},\mathbf{y})=0$ indicates the system will return the output $\mathbf{y}$.
@@
  • Query function $a:\mathcal{X}\rightarrow[0,1]$. $a(\mathbf{x})$ represents whether an input $\mathbf{x}$ can be answered.
@@
The refusal function $r$ determines whether the LLM should provide output $\mathbf{y}$ in response to input $\mathbf{x}$ as:
$r(\mathbf{x},\mathbf{y})=\begin{cases}1&\textrm{if }a(\mathbf{x})<\theta_{a}\textrm{ or }c(\mathbf{x},\mathbf{y})<\theta_{c}\\ &\textrm{ or }h(\mathbf{x},\mathbf{y})<\theta_{h}\\ 0&\textrm{otherwise}\end{cases}$
where $\theta_{a}$, $\theta_{c}$, and $\theta_{h}$ are score thresholds for each of the three functions.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Abstention methods in LLMs, ver, color=black!100, fill=black!15, text=black [ Pretraining, color=yellow!100, fill=yellow!100, text width=4.0em, text=black [ Data augmentation, color=yellow!100, fill=yellow!60, text width=6.8em, text=black [ Neeman et al. (2023), text width=37.5em ] ] ] [ Alignment, color=cyan!100, fill=cyan!100, text width=4.0em, text=black [ Instruction tuning, color=cyan!100, fill=cyan!60, text width=6.8em, text=black [ Yang et al. (2023); Zhang et al. (2024a); Lyu et al. (2024); Wolfe et al. (2024); Feng et al. (2024b); Wang et al. (2024d)
Zhang et al. (2023b); Brahman et al. (2024); Bianchi et al. (2024); Wallace et al. (2024); Varshney et al. (2023), text width=37.5em ] ] [ Learning from
@@
preferences, color=cyan!100, fill=cyan!60, text width=6.8em, text=black [ Zhang et al. (2024b); Cheng et al. (2024); Liang et al. (2024); Guo et al. (2024); Shi et al. (2024); Bai et al. (2022)
@@
Lin et al. (2024a); Dai et al. (2024); Touvron et al. (2023); Brahman et al. (2024); Sun et al. (2024); Kim et al. (2024a), text width=37.5em ] ] ] [ Inference, color=lightgreen!200, fill=lightgreen!200, text width=4.0em, text=black [ Input-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Query processing, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Cole et al. (2023); Qi et al. (2021); Hu et al. (2024); Sun et al. (2023); Xi et al. (2023); Jain et al. (2023)
@@
Shao et al. (2021); Kumar et al. (2024); Dinan et al. (2019), text width=33em ] ] ] [ In-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Probing LLM’s
@@
inner state, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Kadavath et al. (2022); Azaria and Mitchell (2023); Wang et al. (2024a); Liu et al. (2020); Chen et al. (2024)
Kamath et al. (2020); Slobodkin et al. (2023); Bhardwaj et al. (2024), text width=33em ] ] [ Uncertainty
@@
estimation, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Lin et al. (2022); Tian et al. (2023); Tomani et al. (2024); Kadavath et al. (2022); Duan et al. (2023)
Xiong et al. (2024); Shrivastava et al. (2023); Achiam et al. (2023); Cole et al. (2023); Kuhn et al. (2023)
@@
Ren et al. (2023); Zhou et al. (2024b), text width=33em ] ] [ Calibration-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Zhao et al. (2022); Xiao et al. (2022); Jiang et al. (2021); Hou et al. (2023); Varshney et al. (2022)
Clark et al. (2020b); Mielke et al. (2022); Lu et al. (2022); Wen et al. (2020); Zablotskaia et al. (2023)
@@
Desai and Durrett (2020); Stengel-Eskin et al. (2024), text width=33em ] ] [ Consistency-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Slobodkin et al. (2023); Xiong et al. (2024); Zhao et al. (2024); Lin et al. (2024b); Ji et al. (2024)
@@
Cao et al. (2023); Cole et al. (2023); Yuan et al. (2024); Chen et al. (2024); Robey et al. (2023), text width=33em ] ] [ Prompting-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Lin et al. (2024b); Slobodkin et al. (2023); Wen et al. (2024); Yang et al. (2023); Cheng et al. (2024)
Zhang et al. (2023b, 2024c); Zheng et al. (2024); Zhou et al. (2024a, c); Ren et al. (2023)
@@
Mo et al. (2024); Pisano et al. (2023); Wei et al. (2024); Xie et al. (2023); Varshney et al. (2023), text width=33em ] ] ] [ Output-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Self evaluation, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Phute et al. (2024); Kadavath et al. (2022); Varshney et al. (2023); Chen et al. (2023b); Ren et al. (2023)
@@
Wang et al. (2024e); Feng et al. (2024a), text width=33em ] ] [ LLM collaboration, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Mielke et al. (2022); Wang et al. (2024b); Pisano et al. (2023); Chen et al. (2023a); Feng et al. (2024b)
Zeng et al. (2024), text width=33em ] ] ] ] ]
@@ 3 Abstention Methodology
We classify abstention methodology (Fig. 2) from an LLM-centered perspective, based on when they are applied in the LLM lifecycle: pretraining, alignment, or inference. Below, we summarize prior work and provide ideas for future experiments.

### v3

@@ Abstract
Abstention, the refusal of large language models (LLMs) to provide an answer, is increasingly recognized for its potential to mitigate hallucinations and enhance safety in LLM systems. In this survey, we introduce a framework to examine abstention from three perspectives: the query, the model, and human values. We organize the literature on abstention methods, benchmarks, and evaluation metrics using this framework, and discuss merits and limitations of prior work. We further identify and motivate areas for future research, such as whether abstention can be achieved as a meta-capability that transcends specific tasks or domains, and opportunities to optimize abstention abilities in specific contexts. In doing so, we aim to broaden the scope and impact of abstention methodologies in AI systems. [^1] A list of abstention-related papers from this review can be found at https://github.com/chenjux/abstention.
@@ 1 Introduction
Large language models (LLMs) have demonstrated generalization capabilities across NLP tasks such as question answering (QA) (Wei et al., 2022a; Chowdhery et al., 2022), abstractive summarization (Zhang et al., 2023a), and dialogue generation (Yi et al., 2024). But these models are also unreliable, having a tendency to “hallucinate” false information in their responses Ji et al. (2023b), generate overly certain or authoritative responses Zhou et al. (2024b), answer with incomplete information Zhou et al. (2023b), or produce harmful or dangerous responses Anwar et al. (2024). In these situations, the model should ideally abstain: to refuse to answer in the face of uncertainty (Wen et al., 2024; Feng et al., 2024b; Yang et al., 2023).
@@
Current methods to encourage abstention typically rely on calibration techniques, including linguistic calibration (Mielke et al., 2022; Huang et al., 2024b), which aim to accurately and consistently estimate a model’s confidence in its response, then arrange for the model to abstain if the confidence score for a given response falls below some threshold (Varshney et al., 2022; Xiao et al., 2022; Desai and Durrett, 2020). But questions of whether a query is aligned with human values or is answerable at all are difficult to model in terms of model confidence Yang et al. (2023).
Prior work demonstrates the potential of abstention to enhance model safety and reliability in constrained settings Varshney et al. (2023); Wang et al. (2024c); Zhang et al. (2024a). In this survey, we attempt to bring together relevant work studying abstention strategies or leading to abstention behaviors, across the diverse range of scenarios encountered by general-purpose chatbots engaging in open-domain interactions. Our goals are to identify gaps and encourage new methods to achieve abstention. Developing or adapting abstention mechanisms to suit a wide array of tasks will enhance the overall robustness and trustworthiness of LLM interactions.
@@
  • We introduce a framework to analyze abstention capabilities from three perspectives that have typically been considered in isolation—query answerability, the confidence of the model to answer the query, and alignment of query and responses with human values. Our framework helps us identify existing research that is relevant to abstention as well as abstention mechanisms that have been developed in prior work (§2).
  • We conduct a detailed survey of existing abstention methods(§3) as well as evaluation benchmarks and metrics (§4), aiding researchers in selecting appropriate strategies. For each class of methods, we identify opportunities for further research to advance the field.
  • We discuss other considerations and under-explored aspects (§5) of abstention, highlighting pitfalls and promising future directions. We encourage researchers to develop more robust model abstention mechanisms and demonstrate their effectiveness in real-world applications.
@@ 2 Abstention in LLMs
We define abstention as the refusal to answer a query. When a model fully abstains, it may begin a response with “I don’t know” or refuse to answer in another way. In reality, abstention encompasses a spectrum of behaviors  Röttger et al. (2024a), e.g., expressing uncertainty, providing conflicting conclusions, or refusing due to potential harm are all forms of abstention. Partial abstention may involve both answering and abstention, such as self-contradictory responses, e.g., “I can’t answer the question, but I suppose the answer might be…” We do not consider ignoring and/or reframing the question as abstention; but rather as failure modes of LLMs in following instructions (Röttger et al., 2024a; Varshney et al., 2023).
For the abstention expression—the words a model uses to convey that it has abstained—we adopt the definition of five major types of expressions from prior work (Varshney et al., 2023; Wang et al., 2024c), indicating that the model (i) cannot assist; (ii) refutes the query; (iii) provides multiple perspectives without expressing preference; (iv) perceives risk associated with the query and answers cautiously with a disclaimer; and (v) refuses to offer concrete answers due to the lack of knowledge or certainty. Expressions can be identified through heuristic rules and key word matching (Zou et al., 2023; Wen et al., 2024; Yang et al., 2023), or through model-based or human-based evaluation (§4).
@@ 2.1 Abstention Framework
  • The model knowledge perspective examines the capabilities of the AI model itself, including its design, training, and inherent biases (Ahdritz et al., 2024; Kim and Thorne, 2024; Hestness et al., 2017; Hoffmann et al., 2022; Kaplan et al., 2020; Cao, 2024). For any given query, the system should abstain if the model is insufficiently confident about the correctness of output or has a high probability of returning an incorrect output.
@@ 2.2 Problem Formulation
To formalize our definition of abstention: consider an LLM $f:\mathcal{X}\rightarrow\mathcal{Y}$. When given a prompt $\mathbf{x}\in\mathcal{X}$, $f$ generates a response $\mathbf{y}\in\mathcal{Y}$. We model refusal to answer (abstention) as a function $r:\mathcal{X,Y}\rightarrow[0,1]$ where $r(\mathbf{x},\mathbf{y})=1$ indicates the system will fully abstain from answering, $r(\mathbf{x},\mathbf{y})=0$ indicates the system will return the output $\mathbf{y}$, and intermediate values represent partial abstention.
@@
  • Query function $a:\mathcal{X}\rightarrow[0,1]$. $a(\mathbf{x})$ represents the degree to which an input $\mathbf{x}$ can be answered.
@@
The refusal function $r$ determines whether the LLM should abstain from responding to input $\mathbf{x}$ as:
$r(\mathbf{x},\mathbf{y})=\begin{cases}1,\textrm{if any of }a(\mathbf{x});c(\mathbf{x},\mathbf{y});h(\mathbf{x},\mathbf{y})=0\\ M(a(\mathbf{x}),c(\mathbf{x},\mathbf{y}),h(\mathbf{x},\mathbf{y})),\textrm{otherwise}\\ 0,\textrm{if all of }a(\mathbf{x});c(\mathbf{x},\mathbf{y});h(\mathbf{x},\mathbf{y})=1\end{cases}$
The function $M$ acts as a connector between the three perspectives, defining how the system integrates their individual outputs into a unified decision. The design of $M$ depends on the system designer and will vary based on the application; examples include weighted averaging, logical operations, or custom thresholds. Some existing systems (Varshney et al., 2022; Cole et al., 2023) use threshold-based mechanisms to convert partial abstention behaviors into binary decisions, such as full compliance or full refusal.
Our framework allows nuanced handling of abstention, by combining confidence from all three perspectives and enabling partial abstention when appropriate. Under this definition, a system would fully abstain from answering if any of the three perspectives indicates full abstention. In all other cases, a system would partially abstain, balancing between providing an answer and withholding information based on indications from the three perspectives.
2.3 Inclusion in this Survey
We identify and survey prior work that falls under any of the three perspectives of our abstention framework. In §3, we organize abstention methodology from an LLM-centered perspective, based on when each method is applied in the LLM lifecycle: pretraining, alignment, or inference. This organization is chosen for ease of comparison of experimental settings. Each subsection within §3 is further organized by the three perspectives. Following, §4 describes evaluation benchmarks and metrics that have been used or introduced across the surveyed prior work. At the end of each subsection, in blue boxes, we summarize main takeaways and provide suggested directions for future work. In §5, we summarize notable threads of research that are not easily classified as method or evaluation.
@@ 2.3 Inclusion in this Survey
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Abstention methods in LLMs, ver, color=black!100, fill=black!15, text=black [ Pretraining, color=yellow!100, fill=yellow!100, text width=4.0em, text=black [ No prior work, color=yellow!100, fill=yellow!60, text width=6em, text=black ] ] [ Alignment, color=cyan!100, fill=cyan!100, text width=4.0em, text=black [ Supervised Fine-tuning, color=cyan!100, fill=cyan!60, text width=8.7em, text=black [ Neeman et al. (2023); Yang et al. (2023); Zhang et al. (2024a); Feng et al. (2024b); Qi et al. (2024); Lyu et al. (2024)
Cheng et al. (2024); Brahman et al. (2024); Gekhman et al. (2024); Lin et al. (2024a); Kang et al. (2024)
Wolfe et al. (2024); Hu et al. (2022); Zhao et al. (2022); Jiang et al. (2021); Lin et al. (2022); Mielke et al. (2022)
@@
Bianchi et al. (2024); Wallace et al. (2024); Varshney et al. (2023); Zhang et al. (2023b); Wang et al. (2024d), text width=46.7em ] ] [ Learning from preferences, color=cyan!100, fill=cyan!60, text width=10em, text=black [ Zhang et al. (2024c); Cheng et al. (2024); Liang et al. (2024); Guo et al. (2024); Shi et al. (2024); Bai et al. (2022)
Kang et al. (2024); Stengel-Eskin et al. (2024); Lin et al. (2024a); Dai et al. (2024); Touvron et al. (2023)
@@
Brahman et al. (2024); Sun et al. (2024), text width=45.5em ] ] ] [ Inference, color=lightgreen!200, fill=lightgreen!200, text width=4.0em, text=black [ Input-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Query processing, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Cole et al. (2023); Min et al. (2020); Qi et al. (2021); Hu et al. (2024); Sun et al. (2023); Xi et al. (2023)
@@
Jain et al. (2023); Shao et al. (2021); Kumar et al. (2024); Dinan et al. (2019), text width=40.1em ] ] ] [ In-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Probing LLM’s
@@
inner state, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Kadavath et al. (2022); Azaria and Mitchell (2023); Wang et al. (2024a); Liang et al. (2024)
Chen et al. (2024); Kamath et al. (2020); Slobodkin et al. (2023); Bhardwaj et al. (2024), text width=40.1em ] ] [ Uncertainty
@@
estimation, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Lin et al. (2022); Tian et al. (2023); Tomani et al. (2024); Kadavath et al. (2022); Varshney et al. (2022)
Xiong et al. (2024); Shrivastava et al. (2023); Achiam et al. (2023); Cole et al. (2023); Duan et al. (2023)
@@
Kuhn et al. (2023); Ren et al. (2023); Zhou et al. (2024b, 2023a); Jung et al. (2024); Hou et al. (2023), text width=40.1em ] ] [ Calibration-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Zhao et al. (2022); Xiao et al. (2022); Jiang et al. (2021); Hou et al. (2023); Varshney et al. (2022)
@@
Wen et al. (2020); Zablotskaia et al. (2023); Desai and Durrett (2020); Gal and Ghahramani (2016), text width=40.1em ] ] [ Consistency-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Slobodkin et al. (2023); Xiong et al. (2024); Zhao et al. (2024b); Lin et al. (2024b); Ji et al. (2024)
@@
Cao et al. (2023); Cole et al. (2023); Yuan et al. (2024b); Chen et al. (2024); Robey et al. (2023), text width=40.1em ] ] [ Prompting-based, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Lin et al. (2024b); Slobodkin et al. (2023); Wen et al. (2024); Yang et al. (2023); Cheng et al. (2024)
Zhang et al. (2023b, 2024d); Mo et al. (2024); Zhou et al. (2024a, c); Ren et al. (2023); Deng et al. (2024)
@@
Zheng et al. (2024); Pisano et al. (2023); Wei et al. (2024); Xie et al. (2023); Varshney et al. (2023), text width=40.1em ] ] ] [ Output-processing, color=lightgreen!150, fill=lightgreen!150, text width=7em, text=black [ Self evaluation, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Phute et al. (2024); Kadavath et al. (2022); Varshney et al. (2023); Chen et al. (2023b)
@@
Ren et al. (2023); Feng et al. (2024b); Kim et al. (2024a); Wang et al. (2024d), text width=40.1em ] ] [ LLM collaboration, color=lightgreen!100, fill=lightgreen!60, text width=6.8em, text=black [ Feng et al. (2024b); Wang et al. (2024b); Pisano et al. (2023); Chen et al. (2023a); Zeng et al. (2024), text width=40.1em ] ] ] ] ]
@@ 3 Abstention Methodology
We summarize methods introduced in prior work (Fig. 2 organizes these by stages in the LLM lifecycle) and provide ideas for future experiments.